OpenStudy (laddiusmaximus):

find an equation of the tangent to the curve at the given point y=sin^2x (pi/6, 1)

OpenStudy (anonymous):

take the derivative, substitute in your x and y values. that will give you the slope of the line. then find the y value of your point by substituting pi/6 into x in the original equation. you should then have the slope and a point. use point slope formula and determine your tangent line.

OpenStudy (anonymous):

like I said, 1st find the derivative of your equation

OpenStudy (laddiusmaximus):

I tried using the product rule and the chain rule, but its not working.

OpenStudy (laddiusmaximus):

yeah Im trying but im not doing it right

OpenStudy (anonymous):

You want to find the derivative of \[y=\sin^{2x}\], correct?

OpenStudy (laddiusmaximus):

which is cos^2x

OpenStudy (anonymous):

Ok, so what is the meaning of a derivative?

OpenStudy (laddiusmaximus):

and equation of the tangent line

OpenStudy (anonymous):

nope, the derivative is the slope of the tangent line

OpenStudy (laddiusmaximus):

ok so now what? I just plug in pi/6?

OpenStudy (anonymous):

Yes, plug pi/6 in to the derivative as x, the result will be the slope of the tangent line

OpenStudy (laddiusmaximus):

but what about 1?

OpenStudy (laddiusmaximus):

so I have 4cos^2(pi/6)?

zepdrix (zepdrix):

\[\large y=\sin^2x=(sinx)^2\]\[\large y'=2(sinx)(sinx)'=2(sinx)(cosx)\] Did you find the derivative ok?

zepdrix (zepdrix):

When we take derivatives we work from the outside in. We take the derivative of the OUTERMOST function, then apply the chain rule, multiplying by the derivative of the inside. The problem with trig functions is the notation. Sometimes it's tough to recognize what the outermost function is because of the placement of the exponent.

OpenStudy (laddiusmaximus):

so 4sin^2x becomes 4sinx cosx ?

zepdrix (zepdrix):

Don't forget that factor of 2 in front! 4sin^2x would become 4[2(sinx)(cosx)]=8sinxcosx

OpenStudy (anonymous):

Oh my goodness, zepdrix is correct. I did not check your derivation.

OpenStudy (anonymous):

I apologize for that.

zepdrix (zepdrix):

Confused lad? D':

OpenStudy (laddiusmaximus):

so i plug pi/6 and 1 in seperately?

OpenStudy (laddiusmaximus):

okay I get the chain rule. I just didnt see it when I looked at the problem. so I have the derivative now what?

zepdrix (zepdrix):

\[\large \sin^2(\frac{ \pi }{ 6 })\neq 1\] Hmm I'm not sure where that (pi/6, 1) is coming from. It's not an ordered pair. Oh oh oh oh. Is your problem suppose to be this? \[\large y=4\sin^2x\]

zepdrix (zepdrix):

Remember what a tangent line looks like? It touches our curve at a particular point, and has the same slope as that curve at the given point. Our job is to find an equation for that straight line. We'll put it in the form y=mx+b.

OpenStudy (laddiusmaximus):

yeah crap sorry its y=4sin^2x

zepdrix (zepdrix):